OpenStudy (anonymous):

GIVING MEDALS FOR BEST ANSWER!!!!!!! Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (akashdeepdeb):

Do you know about the relation of slopes of 2 perpendicular lines? m1.m2 = -1 Know that?

OpenStudy (akashdeepdeb):

So written in Slope Intercept form [y=mx+b] y = -4x -2 [The slope m1 = -4] Then the slope of the line perpendicular to this line would be m2 where m1.m2 = -1 m2 = 1/4 NOTE: m1 is a **NEGATIVE RECIPROCAL** of m2. So now you have slope = 1/4 a point = (-16,-11) Use Point-slope form of a line [(y-y1) = m(x-x1)] to find the equation and then represent it in slope intercept form. :) Understood?

OpenStudy (anonymous):

Nope

OpenStudy (akashdeepdeb):

What did you NOT understand?

OpenStudy (akashdeepdeb):

@Hello_Im_Nick ?

OpenStudy (anonymous):

I will figure it out on my own! Thanks for the help anyway. You deserve a medal :D I'm starting to get a bit more familiar with the subject so thanks anyways! :D

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

YOU LIKE MINECRAFT?! ME TOO!